# Reconstructing Hand-Held Objects in 3D

### v2

@@ Abstract
Objects manipulated by the hand (i.e., manipulanda) are particularly challenging to reconstruct from in-the-wild RGB images or videos. Not only does the hand occlude much of the object, but also the object is often only visible in a small number of image pixels. At the same time, two strong anchors emerge in this setting: (1) estimated 3D hands help disambiguate the location and scale of the object, and (2) the set of manipulanda is small relative to all possible objects. With these insights in mind, we present a scalable paradigm for hand-held object reconstruction that builds on recent breakthroughs in large language/vision models and 3D object datasets. Our model, MCC-Hand-Object (MCC-HO), jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs. Subsequently, we use GPT-4(V) to retrieve a 3D object model that matches the object in the image and rigidly align the model to the network-inferred geometry; we call this alignment Retrieval-Augmented Reconstruction (RAR). Experiments demonstrate that MCC-HO achieves state-of-the-art performance on lab and Internet datasets, and we show how RAR can be used to automatically obtain 3D labels for in-the-wild images of hand-object interactions.
@@ 1 Introduction
“Man is the measure of all things.”
Protagoras
In manipulation, one may philosophize that the hand is the measure of all objects. From infancy, humans use their hands to gather information about the physical world [30, 83, 64, 67]. Through our hands and eyes, we learn to estimate 3D geometry and physical properties as well as affordances that inform the ways in which we manipulate objects [39, 14].
Recovering hand-object interactions in 3D is an important problem in both computer vision and robotics. While there has been significant progress in reconstructing hands [63, 50, 52] and objects [85, 37, 36] separately, joint reconstruction of hand-object interactions presents unique challenges that cannot currently be solved by general reconstruction approaches. Procedural methods such as Structure from Motion (SfM) [72] will struggle because the object is mostly occluded, and thus it is difficult to obtain enough feature points to achieve accurate reconstruction. Similarly, Neural Radiance Fields (NeRFs) [44] are best suited for unoccluded multiview reconstruction of static scenes and objects. Instead, state-of-the-art approaches in hand-object reconstruction have relied on human assistance for data collection [3, 51, 13]. While manually-retrieved 3D object models naturally improve reconstruction quality and can provide better initialization, this paradigm is not scalable. This is because human-driven retrieval typically includes searching online for a suitable object model, verifying that the model matches the object in the image, and providing an estimate of 3D object pose [3]. Rigidly aligning the 3D model to an image is also an underconstrained 3D-to-2D problem, and optimization approaches are prone to issues with local minima.
Figure 1: We present a scalable approach to hand-held object reconstruction that is guided by recognition and retrieval. In the first stage, our transformer-based model (MCC-HO) infers hand-object 3D geometry from a single RGB image (Row 2, visualized as meshes). In the second stage, a template object model is obtained using Retrieval-Augmented Reconstruction (RAR). This template model is then rigidly aligned with the output of MCC-HO using ICP (Row 3).
In this paper, we present a novel approach to reconstructing hands manipulating common objects from monocular RGB image(s) or video. We draw from three key insights:
  1. 1.
    One can recover shape even from single images at test time, if the object being seen has some family resemblance to previously seen 3D objects. For children, the ability to see 3D from motion comes first, followed by binocular stereopsis, and then finally they can guess 3D shape even from a single image [92]. An operationalization of this in computer vision is the MCC system [85] — the model requires 3D information at training time, but at test time a single image is enough.
  2. 2.
    In images of hand-held objects, we expect to see hands. The hands also provide a strong anchor for 3D location and scale; for instance, the hands provide a convenient measurement tool for many commonplace activities (e.g., cooking, crafts). Recent progress in single-image hand reconstruction [52] enables the development of a robust hand-centric approach to object reconstruction.
  3. 3.
    Multiple datasets have now been collected which have videos of hand-object interactions with underlying 3D ground truth (see Section 2), but there has not been a model trained on all available data.
Given an RGB image and estimated 3D hand, we train a transformer-based model, MCC-Hand-Object (MCC-HO), that jointly infers 3D hand and object geometry. This geometry is represented as a neural implicit surface composed of occupancy, color, and hand-object segmentation. Our model architecture is adapted from an existing object reconstruction model, MCC [85], and thus we make use of a learned object prior by fine-tuning an MCC model that is trained on the CO3Dv2 dataset [59].
In conjunction with MCC-HO, we propose Retrieval-Augmented Reconstruction (RAR), a method for automatically retrieving object models using large language/vision models to improve network-inferred geometry. Specifically, we prompt GPT-4(V) [48] to recognize the hand-held object in the image and provide a detailed text description. This description is passed to Genie [42], a text-to-3D generative model, to obtain a 3D object. After the object is retrieved, iterative closest point (ICP) [1] or other point cloud fitting techniques can be used to rigidly align the Genie mesh with the estimated geometry from our network. See Figure 1.
@@
To summarize, this paper presents a paradigm for reconstructing hand-held objects that combines the respective advantages of model-free and model-based approaches. Our method consists of three stages: (1) learning-based hand-object reconstruction, (2) object model retrieval, and (3) rigid alignment of the retrieved model with network-inferred geometry. In the first stage, MCC-HO jointly estimates hand and object geometry from a single RGB image and 3D hand. In the second stage, RAR is used to improve reconstruction quality via recognition and retrieval. The hand-held object is recognized by GPT-4(V) and a 3D object is retrieved using a text-to-3D generative model (e.g., Genie [42]). In the third stage, the retrieved 3D object is rigidly aligned with the network-inferred object geometry. Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [5, 3, 38]. To demonstrate the scalability of our approach, we additionally use RAR to obtain ground truth 3D hands and objects for previously unlabeled images in the 100DOH dataset [66].
@@ 2 Related Work
Hand reconstruction. Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [52, 34, 43, 45, 76, 93, 96, 50, 63] or monocular video [28, 82, 88]. These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [62]. Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [77, 94]. Particularly relevant to our work, [61] considers the challenging case of predicting pose when the hand is interacting with objects, which can lead to large occlusions of the hand and/or object.
Model-based hand-object pose estimation. When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [2, 3, 51, 13, 22, 23, 70, 74, 89, 87, 73, 56]. The estimated hands can then be used to rigidly align the object with images or videos [3, 51] via inverse rendering [41, 58]. However, image-based optimization losses are prone to issues with local minima, as depth/3D information is ambiguous from a single RGB image.
Model-free hand-object reconstruction. Reconstructing hand-held objects without corresponding 3D models poses additional challenges compared to general object reconstruction due to frequent occlusions by the hand and significantly less publicly available 3D data of hand-object interactions. Existing approaches are either tailored to reconstruction from a single image [90, 24, 7, 6, 95, 15, 8], RGB video [27, 91, 19, 53], or RGB-D video [80]. Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved template 3D model.
@@
Figure 2: Given an RGB image and estimated 3D hand, our method reconstructs hand-held objects in three stages. In the first stage, MCC-HO is used to predict hand and object point clouds. In the second stage, a template mesh for the object is obtained using Retrieval-Augmented Reconstruction (RAR). In the third stage, the template object mesh is rigidly aligned to the network-inferred geometry using ICP.
@@ 3 Method Overview
We propose a three-stage approach to reconstructing hand-object interactions that combines feedforward network inference with object model retrieval and subsequent 3D fitting. See Figure 2. Our approach is built on two important insights in the hand-object interaction setting. The presence of a human hand provides grounding in terms of both the 3D location and scale of the object. While monocular hand reconstruction is an ill-posed problem and can be a limiting factor for hand-conditioned object reconstruction approaches, recent progress has enabled accurate hand estimation by scaling up [52]. Thus, as in prior work [90, 8], the estimated object pose and shape is conditioned on the estimated 3D hand. Applying a hand-centric approach has several advantages. Since monocular image/video reconstruction is particularly ill-posed with respect to depth and scale, 3D hands can be used to disambiguate the location and size of the object. Furthermore, the location of the 3D hand provides a coordinate system for estimating object geometry; in the case of rotating, symmetric objects, the hand can disambiguate the pose of the object.
In addition, although the hand introduces significant occlusions of the object, the space of possible manipulanda is much smaller than the space of all objects. Thus, this work focuses on reconstructing categories of common household objects, similar in spirit to how the COCO dataset [35] was constructed and in alignment with object [60] and hand-object [38] datasets. The pervasiveness of such commonplace objects enables deploying large image recognition systems for the task of automatically recognizing the object and retrieving a 3D object model. Since household manipulanda are often quite uniform in shape (e.g., a knife, an 8 o.z. mug), the accuracy of such a retrieval process is much higher than for generic objects, whether through lookup in an object dataset [11, 10] or text-to-3D generation methods [42] trained on such data.
4 Hand-Object Geometry Inference
In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [85] for the specific task of hand-object reconstruction. Notably, two domain shifts from the original MCC model are applied. First, the input depth image is removed and the input RGB-D point cloud is replaced with the 3D hand. Second, the model is trained to infer both hand and object geometry, and thus an extra segmentation output is added to the decoder. We call our model MCC-Hand-Object (MCC-HO).
@@ 4 Hand-Object Geometry Inference
The inputs to MCC-HO are a single RGB image and 3D hand geometry. Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM) [29], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image. Hand geometry is typically parameterized by the MANO hand model [62], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$. The MANO hand skeleton contains 21 joints. At training time, the ground truth 3D hand is used. At test time, HaMeR [52] or any other 3D hand estimation methods can be used.
@@ 4.1 Architecture
with barycentric weights $\alpha$. The points $P_{0}$ are used to define a normalized camera world space, such that the 3D coordinates passed to the network are $P=s(P_{0}-\mu(P_{0}))/\sigma(P_{0})$ and geometry is inferred in this same coordinate system. Importantly, $s$ is a constant scaling factor used to ensure the object geometry can be contained by the bounds of this hand-normalized sampling space. See Figure 3.
Figure 3: The input 3D hand is rasterized and per-pixel 3D points are sampled to be passed to the MCC-HO encoder.
@@
4.2 Implementation Details
The training loss is a combination of occupancy, color, and segmentation:
@@ 4.2 Implementation Details
Our model was trained on a single Quandro RTX 8000 for 3k epochs with a learning rate of $1e^{-5}$ and batch size of 16. We use a cosine schedule and a linear warm-up for the first 5 epochs. At each epoch, one random frame from each sequence is selected. Network parameters that correspond to the MCC architecture [85] were initialized using a pretrained MCC model that was trained on the CO3Dv2 dataset [60]. The weights of the additional linear layer in our decoder that outputs hand-object segmentation are initialized via Xavier initialization.
@@ 5 Retrieval-Augmented Reconstruction
In order to further improve upon the network-inferred object geometry, we use GPT-4(V) [48] to automatically detect and retrieve a 3D object corresponding to the hand-held object in an image. We coin this technique Retrieval-Augmented Reconstruction (RAR) in reference to Retrieval-Augmented Generation in natural language processing [32].
@@
Given an input image (one could also use stacked video frames), we prompt GPT-4(V) to provide a detailed description of the object the hand is holding. This text caption is passed to Genie, a text-to-3D model from Luma [42], to retrieve realistic 3D object geometry and appearance (Genie is also used in [33]). See Figure 4. After the object model is obtained, it is rigidly aligned with our network-inferred object geometry using ICP [1]. In particular, we sample a set of random initial rotations for the retrieved 3D model and select the ICP transformation that results in the lowest alignment error. Genie outputs geometry normalized with respect to a unit cube, so the scale of the 3D model is initialized using the bounding volume of the predicted object point cloud.
Figure 4: Retrieval-Augmented Reconstruction (RAR). Given an input image, we prompt GPT-4(V) [48] to recognize and provide a text description of the hand-held object. The text description is passed to a text-to-3D model (in our case, Genie [42]) to obtain a 3D object.

### v3

@@ Abstract
Objects manipulated by the hand (i.e., manipulanda) are particularly challenging to reconstruct from Internet videos. Not only does the hand occlude much of the object, but also the object is often only visible in a small number of image pixels. At the same time, two strong anchors emerge in this setting: (1) estimated 3D hands help disambiguate the location and scale of the object, and (2) the set of manipulanda is small relative to all possible objects. With these insights in mind, we present a scalable paradigm for hand-held object reconstruction that builds on recent breakthroughs in large language/vision models and 3D object datasets. Given a monocular RGB video, we aim to reconstruct hand-held object geometry in 3D, over time. In order to obtain the best performing single frame model, we first present MCC-Hand-Object (MCC-HO), which jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs. Subsequently, we prompt a text-to-3D generative model using GPT-4(V) to retrieve a 3D object model that matches the object in the image(s); we call this alignment Retrieval-Augmented Reconstruction (RAR). RAR provides unified object geometry across all frames, and the result is rigidly aligned with both the input images and 3D MCC-HO observations in a temporally consistent manner. Experiments demonstrate that our approach achieves state-of-the-art performance on lab and Internet image/video datasets. We make our code and models available on the project website: https://janehwu.github.io/mcc-ho
Figure 1: We present a scalable approach to hand-held object reconstruction from monocular RGB images or videos that is guided by object recognition and retrieval. Results demonstrate that our method is able to generate realistic object geometry that is also faithful to visual observations and consistent across frames. Please see the supplementary materials for evaluation of temporal consistency.
@@ 1 Introduction
This paper addresses the problem of recovering hand-object interactions in 3D from monocular RGB images and videos, focusing in particular on reconstructing hand-held object geometry. Given the tremendous advancements in object recognition and retrieval (e.g., generation), such as GPT-4(V) [51] and text-to-3D models [45, 57, 7, 43], a natural question that arises is whether reconstruction can benefit from leveraging these other branches of computer vision. Particularly for the challenging setting of reconstructing hand-held objects, such an approach may be essential to building up available training data to Internet scale. See Figure 1 and the supplementary videos.
Recovering hand-object interactions in 3D is an important problem in both computer vision and robotics. First, we aim to solve a particular instance of a classic computer vision problem (along with recognition and segmentation): 3D reconstruction. It is also a particularly difficult setting because the hand and object need to be reconstructed from a single RGB image, where the hand is articulated and the object can be small with respect to the image size and partially occluded by the hand. For videos, hand-object trajectories need to additionally be temporally consistent. Second, in robotics, particularly manipulation, there is a great need for big data. Unlike other areas such as natural language processing and computer vision, where big data is already available on the Internet, large-scale trajectories of robots manipulating objects are not readily available. There has been a longstanding aspiration to leverage trajectories of human hands interacting with objects [66]; in terms of videos, this data is readily available on the Internet. However, the significant challenge is how to extract 3D hand-object trajectories from those videos. To this end, our objective is to develop an automated technique for taking raw Internet video and extracting 3D hand-object trajectories in a model-free manner. Note that this version of the problem is distinct from model-based approaches that assume a known object shape [55, 100] or use multiview methods to reconstruct the object in 3D in a separate step [13, 31].
@@
Our method consists of three stages: (1) learning-based hand-object reconstruction, (2) object model retrieval, and (3) rigid alignment over time of the retrieved model with network-inferred geometry. In the first stage, our transformer-based model, MCC-Hand-Object (MCC-HO), estimates hand-held object geometry from each RGB image and 3D hand. In the second stage, Retrieval-Augmented Reconstruction (RAR) is used to improve reconstruction quality via recognition and retrieval. Specifically, the hand-held object is recognized by GPT-4(V) and a 3D object is retrieved using a text-to-3D generative model (e.g., Genie [45]). In the third stage, the retrieved 3D object is rigidly aligned with the input image or video using a combination of DINOv2 [52] and inferred 3D cues from MCC-HO. Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [4, 2, 41]. To demonstrate the scalability of our approach, we additionally obtain pseudo ground truth 3D hands and objects for previously unlabeled videos in the 100DOH dataset [71].
Figure 2: Given an RGB video and estimated 3D hands, our method reconstructs hand-held object trajectories in three stages. In the first stage, MCC-HO is used to predict hand and object point clouds. In the second stage, a 3D model for the object is obtained using Retrieval-Augmented Reconstruction (RAR). In the third stage, the 3D object model is rigidly aligned with DINOv2 [52] image features and network-inferred geometry in a temporally consistent manner.
@@ 2 Related Work
Hand reconstruction. Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [56, 37, 46, 48, 82, 98, 101, 53, 69, 59] or monocular video [30, 87, 92]. These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [68]. Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [81, 97]. We use HaMeR [56] in this work as it has achieved state-of-the-art performance on in-the-wild images.
Model-based hand-object pose estimation. When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [1, 2, 55, 14, 24, 25, 74, 78, 93, 91, 77, 62]. The estimated hands can then be used to rigidly align the object with images or videos [2, 55, 100] via inverse rendering [44, 64]. Due to the specific challenges arising from hand-object interactions, e.g. highly occluded geometry, optimization-based approaches are considered the current state-of-the-art in this area. RHO [2] studied the problem of recovering hand-object pose from in-the-wild images, whereas prior work was centered around datasets captured in lab settings. RHOV [55] extended RHO to video data by adding, among other terms, temporal and optical flow losses. Aiming to add more physics-inspired constraints, HandyPriors [100] leverages a differentiable contact module to minimize relative sliding between frames. G-HOP [96] guides optimization with a generative hand-object prior.
Model-free hand-object reconstruction. Reconstructing hand-held objects without corresponding 3D models poses additional challenges compared to general object reconstruction due to frequent occlusions by the hand and significantly less publicly available 3D data of hand-object interactions. Existing approaches are either tailored to reconstruction from a single image [94, 23, 5, 6, 99, 16, 8, 58], RGB video [28, 95, 21, 96], or RGB-D video [85]. Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved 3D object model.
@@ 3 Method Overview
Our approach to reconstructing hand-object interactions from monocular RGB video combines feedforward network inference (MCC-HO) with object model retrieval (RAR) and subsequent rigid alignment. See Figure 2. We build on the insight that the presence of a human hand provides grounding in terms of both the 3D location and scale of the object. Thus, as in prior work [94, 8], the estimated object pose and shape is conditioned on the estimated 3D hand. Applying a hand-centric approach has several advantages. Since monocular image/video reconstruction is particularly ill-posed with respect to depth and scale, 3D hands can be used to disambiguate the location and size of the object. Furthermore, the location of the 3D hand provides a coordinate system for estimating object geometry; in the case of rotating, symmetric objects, the hand can disambiguate the pose of the object.
Given a monocular RGB image or video, our method reconstructs hand-held objects in three stages. First, our transformer-based model, MCC-HO, estimates 3D hand-object geometry from single images (Section 4). Simultaneously, a 3D object model is automatically synthesized by a text-to-3D generative model [45] through a technique we coin Retrieval-Augmented Reconstruction (Section 5). Finally, the 3D object model is rigidly aligned in a temporally consistent manner (Section 6).
4 Estimating Hand-Object Geometry
In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [89] for the specific task of hand-object reconstruction. Our model, MCC-Hand-Object (MCC-HO), has to deal with two major challenges. First, MCC assumes RGB-D images as input, whereas our model needs to work with RGB input, since our ultimate goal is to reconstruct from Internet videos. Second, we have to find a way to cope with the significant domain shift between the training datasets. MCC was trained on the CO3D dataset [67], which consists of clean, 360-degree videos where each object is unoccluded and placed on a solid surface. For MCC-HO, our training data is a relatively small collection of in-the-wild and lab datasets of hand-object interactions, where the hand-held object is at least partially occluded in the vast majority of images.
@@ 4 Estimating Hand-Object Geometry
The inputs to MCC-HO are a single RGB image and 3D hand geometry. Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM/SAM2) [32, 65], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image. Hand geometry is parameterized by the MANO hand model [68], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$. The MANO hand skeleton contains 21 joints. At training time, the ground truth 3D hand is used. At test time, HaMeR [56] is used to infer the hand geometry.
@@ 4.1 Architecture
with barycentric weights $\alpha$. The points $P_{0}$ are used to define a normalized camera world space, such that the 3D coordinates passed to the network are $P=s(P_{0}-\mu(P_{0}))/\sigma(P_{0})$ and geometry is inferred in this same coordinate system. Importantly, $s$ is a constant scaling factor used to ensure the object geometry can be contained by the bounds of this hand-normalized sampling space.
@@
Losses. The training loss is a combination of occupancy, color, and segmentation:
@@ 5 Retrieval-Augmented Reconstruction
Like any other deep learning model, how well MCC-HO is expected to perform is related to the amount of training data available, e.g., the size of available hand-object datasets. These are currently all quite small, orders of magnitude smaller than the “big data” on which computer vision foundation models have been trained. How might we find a way to exploit these other models to our end? Specifically, recognition of objects is very well developed [51], and large collections of 3D object models [10, 11] exist and have facilitated advancements in 3D generative modeling [45, 57, 7, 43]. Thus, in order to further improve upon network-inferred object geometry, we exploit these “big data” models by using GPT-4(V) [51] and a text-to-3D generative model [45] to automatically detect and “retrieve” a 3D object corresponding to the hand-held object in an image. We call this technique Retrieval-Augmented Reconstruction (RAR), somewhat in analogy to Retrieval-Augmented Generation in natural language processing [35].
@@
Given a representative frame of the video, we prompt GPT-4(V) to provide a detailed description of the object the hand is holding. This text caption is passed to Genie, a text-to-3D model from Luma [45], to retrieve realistic 3D object geometry and appearance (Genie is also used in [36]). Since Genie outputs geometry normalized to fit inside a unit cube, the scale of the 3D model must be adjusted with respect to the input 3D hand.
5.1 Rigid Alignment
After a 3D object model is retrieved, it is rigidly aligned with the input visual data and MCC-HO predictions. We begin by considering the single image case. First, the object scale is estimated using the predicted MCC-HO object point cloud, which is a reasonable approximation because the prediction is grounded by the input 3D hand. If the MCC-HO point cloud is $X\in\mathbb{R}^{N\times 3}$ and a sampled point cloud of the retrieved object model is $Y\in\mathbb{R}^{N\times 3}$, then the scale is defined as $s=\sqrt{\lambda_{X}^{0}}/\sqrt{\lambda_{Y}^{0}}$ where $\lambda_{X}^{0}$ is the largest eigenvalue of $X^{T}X$ and $\lambda_{Y}^{0}$ is similarly defined for $Y$.
The remaining parameters to be optimized are object rotation $R$ and translation $T$. In order to (1) guarantee a globally optimal solution while (2) maintaining fixed compute time, we discretize the space of all possible parameter values to construct finite state spaces $\{\mathcal{R},\mathcal{T}\}$. The 3D rotation group, SO(3), is discretized using the unit cube subdivision strategy proposed in [33]. Translation is defined as an offset from the mean location of the MCC-HO point cloud $\mu(X)$, and the space is discretized by constructing a small voxel grid centered at $\mu(X)$. In order to avoid intractably large state space sizes (e.g., $\mathbb{R}^{3}\times\mathbb{R}^{3}$), we first optimize for rotation and then translation. For each parameter, we select the state that results in the lowest error using the sum of the following two metrics.
MCC-HO Chamfer Distance. The Chamfer Distance (CD) between the MCC-HO point cloud and a transformed object model is the usual 3D evaluation metric as defined in Equation 8. Note that when optimizing for rotation, the object model (centered at the origin) is first rotated and then translated to the mean location of the MCC-HO point cloud.
DINOv2 PCA Similarity. The MCC-HO point cloud gives us a reasonable estimate for object pose, but using this signal alone may not result in poses that align well with the input image (particularly for geometry with symmetries and in-the-wild objects). Thus, we additionally use DINOv2 [52] features to determine visual similarities between a transformed object model and the input image. For each rotation/translation state, the transformed object mesh is rendered using Pytorch3D [64]. A PCA basis is constructed using the DINOv2 features of the input image, and the first three PCA components are computed for both the rendered features $\mathcal{F}_{j}$ and input image features $\mathcal{F}_{0}$. We compute the cosine similarity between the two as
${\textstyle E_{DINO}=1-\frac{1}{2}\left(\frac{\text{PCA}(\mathcal{F}_{j})\cdot\text{PCA}(\mathcal{F}_{0})}{\max(||\text{PCA}(\mathcal{F}_{j})||_{2}\cdot||\text{PCA}(\mathcal{F}_{0})||_{2}),\epsilon)}+1\right)}$ (6)
where $\mathcal{F}_{j}$ and $\mathcal{F}_{0}$ are both masked by the input image object silhouette. When ground truth object masks are not available, we use SAM 2 [65]. See Figure 3 for rotation results using example images from 100DOH [71].
Figure 3: We compute a PCA basis of DINOv2 features using all frames masked by the object silhouettes (one frame and its first three PCA components are shown on the left side of each example). The first three components of this basis are used to determine the maximum likelihood Genie object rotation for each frame (shown on the right side).
6 From Images to Videos
In the previous section, we showed how given a single image, we select the rotation and translation of the 3D object model that minimizes error metrics based on CD and DINOv2 similarity. Now, if we have as input a video stream, we can further exploit the additional cue of temporal smoothness—we expect adjacent frames to have rotation and translation states that are close to each other.
Given a sequence of video frames with timesteps $t$, we jointly optimize for per-frame rotations $R(t)$ and subsequently per-frame translations $T(t)$. The typical approach for temporal alignment is to use gradient-based optimization  [55, 95], which is prone to issues with local minima and has unbounded computation time. Aiming for robustness and efficiency, we propose using the Viterbi algorithm [80, 15, 29] to solve for a global solution to rigid object alignment across all frames of a video sequence. The Viterbi algorithm is a dynamic programming approach to obtaining the maximum a posteriori probability of the most likely sequence of hidden states $Q=q_{1},q_{2},\ldots,q_{N}$ that results in a sequence of observed events $O=o_{1},o_{2},\ldots,o_{N}$. The hidden states $q_{t}$ can take on values from a finite set of parameters $\mathcal{Q}$ (where in our case, $\mathcal{Q}=\mathcal{R}$ or $\mathcal{Q}=\mathcal{T}$). Following the single image case, the observations $o_{t}$ are a combination of the MCC-HO inferred object point clouds and the DINOv2 image features.
Each cell of the Viterbi trellis, $v_{t}(j)$, is the probability of ending up in state $j$ after seeing the first $t$ observations and passing through the most probable state sequence $q_{1},\ldots,q_{t-1}$. Given the probability of being in every state at time $t-1$, the Viterbi probabilities are computed via:
$v_{t}(j)=\max_{i=1}^{|\mathcal{Q}|}\hskip 2.84526ptv_{t-1}(i)a_{ij}b_{j}(o_{t}),$ (7)
where $a_{ij}$ is the transition probability from the previous state $q_{i}$ to the current state $q_{j}$, and $b_{j}(o_{t})$ is the state observation likelihood given the current state $j$. In particular, $a_{ij}$ is Rodrigues’ rotation error when optimizing for rotations and Euclidean distance when optimizing for translations. The emission cost $b_{j}(o_{t})$ is the sum of the two metrics defined in the single image discussion, e.g., CD with respect to the MCC-HO point clouds and DINOv2 PCA similarity with respect to the input images.